OpenStudy (anonymous):

has anyone done this test? Lesson 9: Solving Equations Unit Test Algebra 1 A Unit 2: Solving Equations

OpenStudy (anonymous):

Is this for Flvs?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

CA

OpenStudy (anonymous):

Hmmmm. Whats the question? Maybe i can help:)

OpenStudy (anonymous):

-5=x+3

OpenStudy (anonymous):

haha wrong one lol

OpenStudy (anonymous):

-8 is for that

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

that was hard lmao but here is the real one.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

-2=11 over 17z

OpenStudy (anonymous):

17z

OpenStudy (anonymous):

z=-11/34

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

can you explain how you solved?

OpenStudy (anonymous):

and your answer is not one of the choices. lol

OpenStudy (anonymous):

−2=11/17z −(17z)2=11 −17z⋅2=11 −17⋅2z=11 −34z=11 34z/34=−11/34 z=−11/34

OpenStudy (anonymous):

oh okay thatnks!

OpenStudy (anonymous):

Thank you. Do you have anymore?

OpenStudy (anonymous):

yeah but i dont want to cheat on the whole test.

OpenStudy (anonymous):

Come on why not, i do it! lol